solve the linear equation 2x+3y=1,5x+7y=3
Answer by mananth(16946)

2 x + 3 y = 1 .............1
Total value
5 x + 7 y = 3 .............2
Eliminate y
multiply (1)by -7
Multiply (2) by 3
-14 x -21 y = -7
15 x + 21 y = 9
Add the two equations
1 x = 2
/ 1
x = 2
plug value of x in (1)
2 x + 3 y = 1
4 + 3 y = 1
3 y = 1 -4
3 y = -3
y = -1